The magnetization within a bar of some metal alloy is 1.2à ƒ —106 A/m when the H field is 200 A/m.The permeability à ¢ € ²Ãƒ Ž ¼Ãƒ ¢ € ² can be estimated to be:

A) 5.82à ƒ —10à ¢ ˆ ’4H/m
B) 7.54à ƒ —10à ¢ ˆ ’3H/m
C) 6.72à ƒ —10à ¢ ˆ ’4H/m
D) 3.64à ƒ —10à ¢ ˆ ’3H/m
1
Expert's answer
2021-10-13T09:08:28-0400

1) Determine the magnetic susceptibility:


"\chi=\frac MH=\frac{1.2·10^6}{200}=6000."

2) Hence, determine the permeability of the material:


"\mu=\mu_0(1+\chi)=7.54·10^{-3}\text{ H/m}."